Accrual Based Accounting – Adjusting Journal Entries (AJEs): Say Something, Inc. purchases equipment on 5/1/2019, paying $4,500 in cash. The equipment has a 4-year useful life, $0 salvage value, and Say Something uses the ‘straight line depreciation’ method to allocate the cost of the equipment evenly over its useful life.

  • Record the journal entry for the original purchase of equipment on May 1st, 2019.

  • Record the adjusting entry to recognize Depreciation Expense on December 31st, 2019. Assume Say Something uses an annual accounting period which ends on December 31st, 2019 and adjusting entries are only made at the end of the annual accounting period on 12/31 (i.e. assume no adjusting entries have been recorded yet).

  • Assume the balance of Accumulated Depreciation-Equipment account on April 30th, 2019 was $0 and Say Something, Inc. has no other depreciable assets. What will the Net Book Value of the equipment be on December 31, 2019.

  • Assuming we make the next adjusting entry related to this equipment at the end of the 2020 annual accounting period on 12/31/2020 (again, assume the firm did not purchase any other depreciable assets in 2020).  

    • How much Depreciation Expense would Say Something record in 2020?

    • What will the Net Book Value of the equipment be on December 31, 2020?

  • What if we did not make the adjusting entry on 12/31/19 to recognize depreciation expense? Indicate by how much 2019’s 1) assets, 2) liabilities, 3) revenues, 4) expenses, 5) net income, 6) retained earnings, and 7) SHE would be either under- or overstated if this adjusting entry were not recorded. If no effect, write ‘no effect.
Answer #1

1) Journal entry of original purchase:

Date

Acc Titles

Debit $

Credit $

5/1/2019

Equipment

4500

Cash

4500

(cash purchase of equipment)

2) Adjusting at 2019 end:

12/31/2019

Depreciation expense

750

(4500/4)*(8/12)

Acc Dep.- Equipment

750

(booking of Prorata depreciation on equipment)

Net Book Value of the

equipment on 12/31/2019

would be = 4500-750 = 3750

3) Adjusting at 2020 end:

12/31/2020

Depreciation expense

1125

(4500/4)

Acc Dep.- Equipment

1125

(booking of depreciation on equipment)

Net Book Value of the

equipment on 12/31/2020

would be = 4500-750-1125 = 2625

4) If no adjusting entry been made

on 12/31/2019, then

Assets

Overstated

Liabilities

Understated

Revenues

No effect

Expenses

Understated

Net Income

Overstated

Retained Earnings

Overstated

SH Equity Overstated